Date:
Friday, July 16, 2021
Telescope:
CTIO 1.5-m
Submitter:
Roberto Aviles
Submitter Email:
Observer Support:
None
Night Assistant:
None
Problems:
00:25 UT fails to start exp on HIP071904: file 1123 ThAr still ON: turned it OFF manually.
09:20 UT While on TYC_7015-00985-1 tel moved East. Renamed file 1181 as 'junk', might be useful. NE wind 7.5 [m/s] and HA ~ 02hh31mm.
Glitch, numbering systems changed, do not know when/why. Fixed. (1182 was 0000, etc.) All the program + T0394287035 completed with the Sun at -9.6 degrees of altitude.
Program:
Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
12.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
12
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-07-16 | 12.0 | 0.0 | 0.0 | 0.0 | 0.0 | 12 |
T0446370039 files 1117, 1118 SN~2.5-3 || || T0306648160 file 1138 S/N ~ 3
T0380836882 files 1141, 42, 43 S/N ~ 3 || || GAI2317-2911_gap S/N ~ 4
T0394287035 files 1189, 90, 91 low S/N ~ 1